# TRAIL: Trust-Aware Client Scheduling for Semi-Decentralized Federated Learning

### v2

@@ Abstract
Due to the sensitivity of data, Federated Learning (FL) is employed to enable distributed machine learning while safeguarding data privacy and accommodating the requirements of various devices. However, in the context of semi-decentralized Federated Learning (SD-FL), clients’ communication and training states are dynamic. This variability arises from local training fluctuations, heterogeneous data distributions, and intermittent client participation. Most existing studies primarily focus on stable client states, neglecting the dynamic challenges present in real-world scenarios. To tackle this issue, we propose a trust-aware client scheduling mechanism (TRAIL) that assesses client states and contributions, enhancing model training efficiency through selective client participation. Our focus is on a semi-decentralized Federated Learning framework where edge servers and clients train a shared global model using unreliable intra-cluster model aggregation and inter-cluster model consensus. First, we propose an adaptive hidden semi-Markov model (AHSMM) to estimate clients’ communication states and contributions. Next, we address a client-server association optimization problem to minimize global training loss. Using convergence analysis, we propose a greedy client scheduling algorithm. Finally, our experiments conducted on real-world datasets demonstrate that TRAIL outperforms state-of-the-art baselines, achieving an improvement of 8.7% in test accuracy and a reduction of 15.3% in training loss.
@@ Introduction
In modern AI systems, local data on end devices often contains sensitive or private information, rendering traditional edge AI training architectures impractical (Zhang et al. 2024; Wu et al. 2023). To address security and privacy concerns while minimizing communication costs, a new distributed machine learning framework called Federated Learning (FL) has emerged (Wu et al. 2023; McMahan et al. 2016). In FL, each client uploads only model parameters, safeguarding their local data. Typically, this process involves coordination with a single edge server, which can result in high communication overhead and potential single points of failure, particularly in environments with numerous end devices (Zhang et al. 2023a; Lu et al. 2022).
@@
This paper investigates semi-decentralized Federated Learning (SD-FL) as a framework to enhance the reliability of model training. As illustrated in Figure 1, we focus on a multi-edge server and multi-client SD-FL framework (Sun et al. 2021b). This architecture employs a two-tier aggregation approach: first, intra-cluster aggregation, where local models are aggregated by their respective servers, followed by inter-cluster consensus, where models from multiple servers are exchanged and collaboratively aggregated to train a shared global model. By distributing computational and communication loads, SD-FL improves both the robustness and scalability of the Federated Learning process. While SD-FL mitigates risks associated with single points of failure, existing research often overlooks the dynamic nature of clients, particularly fluctuations in model contributions and communication quality, which can adversely affect training efficiency (Sun et al. 2021a).
@@
Research has been conducted to address the issue of unreliable clients. In (Sefati and Navimipour 2021), the authors introduce an effective service composition mechanism based on a hidden Markov model (HMM) and ant colony optimization to tackle IoT service composition challenges related to Quality-of-Service parameters, achieving significant improvements in availability, response time, cost, reliability, and energy consumption. In (Ma et al. 2021), the FedClamp algorithm is proposed, which enhances the performance of the global model in Federated Learning environments by utilizing HMM to identify and isolate anomalous nodes. This algorithm is specifically tested on short-term energy forecasting problems. The authors of (Vono et al. 2021) present the Quantized Langevin Stochastic Dynamics (QLSD) algorithm, which employs Markov Chain Monte Carlo methods to improve dynamic prediction capabilities in Federated Learning while addressing challenges related to privacy, communication overhead, and statistical heterogeneity. Additionally, (Wang et al. 2024) introduces a trust-Age of Information (AoI) aware joint design scheme (TACS) aimed at enhancing control performance and reliability in wireless communication networks within edge-enabled Industrial Internet of Things (IIoT) systems operating in harsh environments, utilizing a learning-based trust model and scheduling strategy. While these studies explore various dynamic aspects, they do not adequately address the interplay between dynamics and client selection strategies.
To address this gap, we employ a Markov model to predict dynamic changes in client performance and communication quality, allowing us to schedule clients based on these predictions to minimize loss. We propose a joint design scheme that integrates dynamic prediction with client selection, significantly enhancing the control performance and reliability of Federated Learning systems. Extensive experiments and theoretical analyses demonstrate the effectiveness and robustness of our approach under varying client dynamics. The main contributions of this paper are summarized as follows:
@@
  • We propose a unified optimization framework for SD-FL that integrates performance prediction and client scheduling, enhancing model robustness, accelerating convergence speed, and improving overall performance.
@@
  • We introduce an adaptive hidden semi-Markov model (AHSMM) to predict client performance and channel variations to obtain trust levels. This model effectively accounts for both dynamic and static aspects of clients, enabling efficient state predictions for each one.
@@
  • Our experiments on real-world datasets demonstrate that our proposed mechanism outperforms state-of-the-art baselines, achieving an 8.7% increase in test accuracy and a 15.3% reduction in training loss.
@@
In Federated Learning (FL), model training is distributed across multiple clients to protect data privacy and minimize the need for centralized data aggregation. Traditional Federated Learning assumes reliable and frequent communication between clients and the server. However, this assumption often fails in real-world applications, particularly in environments with heterogeneous devices and unstable communication. To address these challenges, researchers have introduced semi-decentralized Federated Learning (SD-FL). This approach combines the benefits of centralized and distributed architectures by enabling direct communication among some clients, thereby reducing the server’s workload and communication costs. SD-FL is better equipped to adapt to dynamic network environments and heterogeneous data distributions, enhancing the system’s robustness and efficiency.
Current research efforts primarily focus on the following areas. Client Selection: Mechanisms are developed to select clients for participation in training, ensuring that chosen clients meet performance criteria, thereby enhancing the overall effectiveness of the model (Lin et al. 2021; Wang et al. 2024; Yemini et al. 2022; Sun et al. 2023). Trust Management: Trust mechanisms are introduced to assess and predict the reliability of clients, ensuring that only reliable clients participate in training. These mechanisms contribute to improved model robustness and performance (Beltrán et al. 2023; Parasnis et al. 2023; Xu et al. 2024; Valdeira et al. 2023).
While there are studies that focus separately on client selection and trust management, existing methods often fail to effectively integrate these aspects. This gap negatively impacts the efficacy and efficiency of SD-FL systems. An integrated approach that addresses both client selection and trust management is essential for enhancing the robustness and performance of SD-FL, especially in diverse and potentially unreliable environments.
@@
Figure 1: system framework.
@@
In this paper, we examine the SD-FL training process across $T$ rounds, involving $S$ edge servers represented by $\mathcal{S}=\{1,2,\cdots,S\}$, and $U$ client devices represented by $\mathcal{U}=\{1,2,\cdots,U\}$. Each round consists of the following steps:
@@
The Adaptive Hidden Semi-Markov Model (AHSMM) extends the traditional Hidden Semi-Markov Model (HSMM) (Dai et al. 2022) by integrating adaptive training to facilitate device training quality diagnostics, communication quality diagnostics, and lifespan prediction and using multi-sensor information, thereby enhancing both modeling and analytical capabilities. The AHSMM model can be described by the parameters $\Sigma=(\pi,A,B,E)$, where: $\pi$ represents the initial state probabilities, $A$ denotes the macro state transition probabilities, $B$ corresponds to the observation probabilities after adaptive training, $E$ represents the state dwell time after adaptive training, encompassing both the existing and remaining dwell times. In addition, similar to HSMM, AHSMM addresses three core problems: evaluation, recognition, and training. To this end, AHSMM defines new forward-backward variables and proposes improved algorithms for forward-backward processes, the Viterbi algorithm (Zhang et al. 2023b), and the Baum-Welch algorithm (Zhang et al. 2023c).
The computational complexity of the Hidden Semi-Markov Model (HSMM) is relatively high. To address this complexity, the Adaptive Hidden Semi-Markov Model (AHSMM) introduces a new forward variable, denoted as $\alpha_{t}(i,e)$. This variable represents the probability of generating the observations $z_{1},z_{2},\ldots,z_{t}$, given that the quality state $i$ has a specific dwell time of $e_{t}(i,e)=e$. In this context, $\varepsilon_{t}$ signifies the current dwell time of the quality state $q_{t}$.
When $\left(q_{t},\varepsilon_{t}\right)$ is assigned the value $(i,e)$, it indicates that the device has remained in its current quality state $i$ up to time $t$. During this period, the state $i$ has accumulated a dwell time of $e$ and is prepared to transition to a different quality state at time $t+$ 1. Therefore, for $11\leq t\leq T-$ 1 and $e\in[1,E]$, we can define the forward variable as follows:
@@
In this context, the variable $\alpha_{T}(i,e)$ is defined as the joint probability of observing the sequence $z_{1},z_{2},\ldots,z_{T}$ while the system is in quality state $i$ over the dwell time interval from $T-e+1$ to $T$. Mathematically, it can be expressed as:
@@
Monitoring a device with multiple sensors can significantly improve quality prediction. Given the inherent differences among sensors, effective data fusion is essential for integrating their information. Consequently, estimating the parameters of the AHSMM becomes necessary. This estimation process utilizes Maximum Likelihood Linear Regression (MLLR) transformations to address the variations across sensors. Simultaneously, a canonical model is trained based on a set of MLLR transformations. Linear transformations are then applied to the mean vectors of the state output and dwell time distributions in the standard model, allowing for the derivation of mean vectors for these distributions.
The formulas are given by:
$\begin{cases}b_{i}\left(z^{(s)}\right)=N\left(Z;\mu_{i}^{(s)}\Sigma_{i}\right)\\ p_{i}(e)=N\left(e;\mu_{e}^{(s)},\sigma_{i}^{2}\right)\quad\\ \mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}\end{cases}.$ (9)
Here, $b_{i}\left(\boldsymbol{z}^{(s)}\right)$ represents the probability density function for the state $i$ based on the observed data $z^{(s)}$ from sensor $s$, modeled as a multivariate normal distribution with mean $\mu_{i}^{(s)}$ and covariance $\Sigma_{i}$. The term $p_{i}(e)$ signifies the probability density function for the dwell time $e$ in state $i$, also following a normal distribution characterized by mean $\mu_{e}^{(s)}$ and variance $\sigma_{i}^{2}$. The mean dwell time $\mu_{e}^{(s)}$ is calculated using the formula $\mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}$, where $\delta^{(s)}$ is a scaling factor, $m_{i}$ represents a parameter related to state $i$, and $\psi^{(s)}$ is a sensor-specific offset.
Joint Estimation of AHSMM Parameters. Let $S$ denote the number of sensors, and let $\boldsymbol{Z}=\left(\boldsymbol{Z}^{(1)},\cdots,\boldsymbol{Z}^{(S)}\right)$ represent the monitoring data, where $\boldsymbol{Z}^{(s)}=\left(z_{1s},\cdots,z_{Ts}\right)$ represents the monitoring data of sensor $s$ with length $T_{s}$.
Here, the parameters are estimated by jointly considering the contributions of all sensors and their respective transformations. The term $\gamma_{t}^{e}(i)$ represents the probability of being in state $i$ with dwell time $d$ at time $t$, $\eta^{(s)}$ and $\xi^{(s)}$ are the transformation matrices and vectors for sensor $s$, and $\Sigma_{i}$ is the covariance matrix for state $i$. This joint estimation process ensures that the model parameters are optimally adjusted for the diverse sensor data.
The MAP estimation of state $q_{t}$ using the AHSMM is calculated in the following way:
@@
In client quality diagnostics and forecasting, information from various sensors often plays distinct roles in decision-making. The AHSMM addresses the challenge of effectively integrating data from multiple sensors.
@@
where $m(i)$ denotes the mean dwell time in state $i$, and $\sigma^{2}(i)$ is the variance of the dwell time in that state. The term $\rho$ serves as a proportionality constant, which adjusts the influence of the variance on the overall residence time.
This formulation captures the idea that the total time spent in a quality state is influenced not only by the average time spent there but also by the variability of that time. A higher variance indicates greater uncertainty in the duration spent in state $i$, which can lead to longer overall residence times. By incorporating both mean and variance, we obtain a more comprehensive view of the dynamics in quality states.
The proportionality constant $\rho$ is defined as:
@@
The reliability function $R(t+e\Delta t)$ represents the probability that the client remains in the current quality state $i$ at time $t+e\Delta t$. Thus:
@@
Based on above equations, $\bar{E}(i,e)$ is expressed as:
@@
Using the previous equations, once the client reaches state $i$ and has a dwell time, the trust level (TL) is determined as:
@@
Considering the above configuration, the model aggregation within the cluster at server $s$ can be described as follows:
$\boldsymbol{w}_{s,t}=\sum_{i\in U_{s}}\frac{n_{i}}{\sum n_{i}}\boldsymbol{w}_{i,t}.$ (17)
@@
$\boldsymbol{g}_{s,t}=\sum_{s=1}^{S}\frac{N_{s}}{\sum N_{s}}\boldsymbol{w}_{s,t}.$ (18)
Furthermore, during the next round of training, each client utilizes the updated local model along with their respective datasets. The local models are trained using the gradient descent mechanism described below. This mechanism allows for distributed learning, where clients can collaboratively train a model while keeping their data localized, thus addressing privacy concerns and reducing bandwidth requirements for data transmission.
@@
Client scheduling aims to determine the optimal client-server association matrix $\boldsymbol{d}$, minimizing the overall training loss. Each element $d_{ij}$ is binary (1 for associated, 0 for not), optimizing data locality, reducing communication overhead, and balancing computational loads. Factors like client data, server capacity, and network conditions influence $\boldsymbol{d}$’s configuration, while adaptive scheduling further enhances performance. The global loss function is:
@@
We rely on current trust levels of client $i$ to tackle these challenges, reduce data loss, and guarantee consistent model updates. $\Theta$ denotes the threshold for the trust level of clients involved in training. Reformulating the optimization problem to incorporate parameters that reflect communication link stability will enhance the modeling of the FL system’s conditions.
We also plan to enhance FL system robustness through redundancy strategies, like multiple communication paths or backup servers, to mitigate risks from unreliable links. Dynamically adjusting client-server associations based on real-time assessments will help maintain optimal performance despite trust fluctuations. This approach maximizes resource utilization and minimizes training time, leading to more robust convergence and broader adoption in real-world applications.
@@
. Setting the learning rate to $\lambda=\frac{1}{L}$ allows us to establish the upper limit of $\mathbb{E}\left(F\left(\boldsymbol{g}_{t+1}\right)-F\left(\boldsymbol{g}^{*}\right)\right)$ in the following manner. The expected difference between the function value at the updated iterate $\boldsymbol{g}_{t+1}$ and the optimal function value $\boldsymbol{g}^{*}$ can be bounded as shown below:
@@
where
$D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L},$ (22)
and
$\begin{cases}B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi,\\ \Psi=\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i,s}<\Theta\right)\right)\right).\end{cases}$ (23)

### v3

@@ Abstract
Due to the sensitivity of data, Federated Learning (FL) is employed to enable distributed machine learning while safeguarding data privacy and accommodating the requirements of various devices. However, in the context of semi-decentralized FL, clients’ communication and training states are dynamic. This variability arises from local training fluctuations, heterogeneous data distributions, and intermittent client participation. Most existing studies primarily focus on stable client states, neglecting the dynamic challenges inherent in real-world scenarios. To tackle this issue, we propose a TRust-Aware clIent scheduLing mechanism called TRAIL, which assesses client states and contributions, enhancing model training efficiency through selective client participation. We focus on a semi-decentralized FL framework where edge servers and clients train a shared global model using unreliable intra-cluster model aggregation and inter-cluster model consensus. First, we propose an adaptive hidden semi-Markov model to estimate clients’ communication states and contributions. Next, we address a client-server association optimization problem to minimize global training loss. Using convergence analysis, we propose a greedy client scheduling algorithm. Finally, our experiments conducted on real-world datasets demonstrate that TRAIL outperforms state-of-the-art baselines, achieving an improvement of 8.7% in test accuracy and a reduction of 15.3% in training loss.
@@ Introduction
In modern AI systems, local data on end devices often contains sensitive or private information, rendering traditional edge AI training architectures impractical (Zhang et al. 2024; Wang et al. 2024a). To address security and privacy concerns while minimizing communication costs, a new distributed machine learning framework called Federated Learning (FL) has emerged (Wu et al. 2023; McMahan et al. 2017; Wang et al. 2024b). In FL, each client uploads only model parameters, safeguarding her local data. Typically, this process involves coordination with a single edge server, which can result in high communication overhead and potential single points of failure, particularly in environments with numerous end devices (Zhang et al. 2023a; Lu et al. 2022).
@@
This paper investigates semi-decentralized FL (SD-FL) as a framework to enhance the reliability of model training. As illustrated in Figure 1, we focus on a multi-edge server and multi-client SD-FL framework (Sun et al. 2021). This framework utilizes a two-tier aggregation approach. The first tier is intra-cluster aggregation, where local models are aggregated by their respective servers. The second tier is inter-cluster consensus, which involves exchanging models from multiple servers and collaboratively aggregating them to train a shared global model. By distributing computational and communication loads, this approach enhances both the robustness and scalability of the FL process. While SD-FL mitigates risks associated with single points of failure, existing research often overlooks the dynamic nature of clients, particularly fluctuations in model contributions and communication quality, which can adversely affect training efficiency (Sun et al. 2023).
@@
Research was conducted to address the issue of unreliable clients. In (Sefati and Navimipour 2021), the authors introduced an effective service composition mechanism based on a hidden Markov model (HMM) and ant colony optimization to tackle IoT service composition challenges related to Quality-of-Service parameters. This approach achieved significant improvements in availability, response time, cost, reliability, and energy consumption. In (Ma et al. 2021), the FedClamp algorithm was proposed, which enhanced the performance of the global model in FL environments by utilizing HMM to identify and isolate anomalous nodes. This algorithm was specifically tested on short-term energy forecasting problems. The authors of (Vono et al. 2021) presented the Quantized Langevin Stochastic Dynamics (QLSD) algorithm, which employed Markov Chain Monte Carlo methods to improve dynamic prediction capabilities in FL while addressing challenges related to privacy, communication overhead, and statistical heterogeneity. Additionally, Wang et al. introduced a trust-Age of Information (AoI) aware joint design scheme (TACS) aimed at enhancing control performance and reliability in wireless communication networks within edge-enabled Industrial Internet of Things (IIoT) systems operating in harsh environments. This scheme utilized a learning-based trust model and scheduling strategy. While these studies explored various dynamic aspects, they did not adequately address the interplay between dynamics and client selection strategies.
To bridge this gap, we propose an adaptive hidden semi-Markov model (AHSMM) to predict dynamic changes in training quality and communication quality. AHSMM enhances standard HMM by explicitly modeling state duration distributions, reducing computational complexity, and adapting to dynamic, multi-parameter environments, making it ideal for complex scenarios. To improve SD-FL systems’ control and reliability, we propose a joint mechanism combining dynamic prediction with client selection. Extensive experiments and analyses validate the effectiveness and robustness of our approach under varying client dynamics. The main contributions of this paper are as follows:
@@
  • We propose a unified optimization mechanism named TRAIL for SD-FL that integrates performance prediction and client scheduling, enhancing model robustness, accelerating convergence speed, and improving overall performance.
@@
  • We introduce an AHSMM to predict client performance and channel variations to obtain trust levels. This model effectively accounts for both dynamic and static aspects of clients, enabling efficient state predictions for each one.
@@
  • Extensive experiments conducted on four real-world datasets (MNIST, EMNIST, CIFAR10, SVHN) demonstrate that our proposed mechanism outperforms state-of-the-art baselines, achieving an 8.7% increase in test accuracy and a 15.3% reduction in training loss.
@@
In FL, model training is distributed across multiple clients to protect data privacy and minimize the need for centralized data aggregation. Traditional FL assumed reliable and frequent communication between clients and the server. However, this assumption often failed in real-world applications, particularly in environments with heterogeneous devices and unstable communication. To address these challenges, researchers introduced SD-FL. This framework combined the benefits of centralized and distributed architectures by enabling direct communication among some clients, thereby reducing the server’s workload and communication costs. SD-FL was better equipped to adapt to dynamic network environments and heterogeneous data distributions, enhancing the system’s robustness and efficiency.
Research efforts primarily focused on the following two areas. (i) Client Selection: Mechanisms were developed to select clients for participation in training, ensuring that chosen clients met performance criteria, thereby enhancing the overall effectiveness of the model (Lin et al. 2021; Wang et al. 2024c; Yemini et al. 2022; Sun et al. 2023). (ii) Trust Management: Trust mechanisms were introduced to assess and predict the reliability of clients, ensuring that only reliable clients participated in training. These mechanisms contributed to improved model robustness and performance (Martínez Beltrán et al. 2023; Parasnis et al. 2023; Xu et al. 2024; Valdeira et al. 2023).
While some studies focus separately on client selection and trust management, existing methods often fail to effectively integrate these aspects, which negatively impacts the efficiency and performance of SD-FL systems. Our work proposes an integrated mechanism that addresses client selection and trust management, which is essential for enhancing the robustness and performance of SD-FL, especially in diverse and potentially unreliable environments.
@@
Figure 1: The SD-FL system framework.
@@
We examine the SD-FL training process across $T$ rounds, involving $S$ edge servers represented by $\mathcal{S}=\{1,2,\cdots,S\}$, and $U$ client devices represented by $\mathcal{U}=\{1,2,\cdots,U\}$. Each round consists of the following steps:
@@
The Adaptive Hidden Semi-Markov Model (AHSMM) extends the traditional Hidden Semi-Markov Model (HSMM) (McDonald, Gales, and Agarwal 2023) into adaptive training using multi-parameter information (i.e., client training accuracy, packet loss), thereby enhancing both modeling and analytical capabilities. The AHSMM model can be described by the parameters $\Sigma=(\pi,A,B,E)$, where: $\pi$ represents the initial state probabilities, $A$ denotes the macro state transition probabilities, $B$ corresponds to the observation probabilities after adaptive training, $E$ represents the state dwell time after adaptive training, encompassing both the existing and remaining dwell times. In addition, similar to HSMM, AHSMM addresses three core problems: evaluation, recognition, and training. To this end, AHSMM defines new forward-backward variables and proposes improved algorithms for forward-backward processes, the Viterbi algorithm (Zhang et al. 2023b), and the Baum-Welch algorithm (Zhang et al. 2023c).
The computational complexity of the Hidden Semi-Markov Model (HSMM) is relatively high. To address this complexity, the Adaptive Hidden Semi-Markov Model (AHSMM) introduces a new forward variable, denoted as $\alpha_{t}(i,e)$. This variable represents the probability of generating the observations $z_{1},z_{2},\ldots,z_{t}$, given that the quality state $i$ has a specific dwell time of $e_{t}(i,e)=e$. In this context, $\varepsilon_{t}$ signifies the current dwell time of the quality state $q_{t}$. When $\left(q_{t},\varepsilon_{t}\right)$ is assigned the value $(i,e)$, it indicates that the device has remained in its current quality state $i$ up to time $t$. During this period, the state $i$ has accumulated a dwell time of $e$ and is prepared to transition to a different quality state at time $t+$ 1. Therefore, for $11\leq t\leq T-$ 1 and $e\in[1,E]$, we can define the forward variable as follows:
@@
The variable $\alpha_{T}(i,e)$ is defined as the joint probability of observing the sequence $z_{1},z_{2},\ldots,z_{T}$ while the system is in quality state $i$ over the dwell time interval from $T-e+1$ to $T$. Mathematically, it can be expressed as:
@@
Monitoring a device with multiple parameters can significantly improve quality prediction. Given the inherent differences among parameters, effective data fusion is essential for integrating their information. Consequently, estimating the parameters of the AHSMM becomes necessary. This estimation process utilizes Maximum Likelihood Linear Regression (MLLR) transformations to address the variations across parameters. Simultaneously, a canonical model is trained based on a set of MLLR transformations. Linear transformations are then applied to the mean vectors of the state output and dwell time distributions in the standard model, allowing for the derivation of mean vectors for these distributions. The formulas are given by:
$\begin{cases}b_{i}\left(z^{(s)}\right)=N\left(Z;\mu_{i}^{(s)}\Sigma_{i}\right),\\ p_{i}(e)=N\left(e;\mu_{e}^{(s)},\sigma_{i}^{2}\right),\quad\\ \mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}.\end{cases}.$ (9)
Here, $b_{i}\left(\boldsymbol{z}^{(s)}\right)$ represents the probability density function for the state $i$ based on the observed data $z^{(s)}$ from parameter $s$, modeled as a multivariate normal distribution with mean $\mu_{i}^{(s)}$ and covariance $\Sigma_{i}$. The term $p_{i}(e)$ signifies the probability density function for the dwell time $e$ in state $i$, also following a normal distribution characterized by mean $\mu_{e}^{(s)}$ and variance $\sigma_{i}^{2}$. The mean dwell time $\mu_{e}^{(s)}$ is calculated using the formula $\mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}$, where $\delta^{(s)}$ is a scaling factor, $m_{i}$ represents a parameter related to state $i$, and $\psi^{(s)}$ is a parameter-specific offset.
Let $S$ denote the number of parameters, and let $\boldsymbol{Z}=(\boldsymbol{Z}^{(1)},\cdots,\boldsymbol{Z}^{(S)})$ represent the monitoring data, where $\boldsymbol{Z}^{(s)}=(z_{1s},\cdots,z_{Ts})$ represents the monitoring data of parameter $s$ with length $T_{s}$. Here, the parameters are estimated by jointly considering the contributions of all parameters and their respective transformations. The term $\gamma_{t}^{e}(i)$ represents the probability of being in state $i$ with dwell time $d$ at time $t$, $\eta^{(s)}$ and $\xi^{(s)}$ are the transformation matrices and vectors for parameter $s$, and $\Sigma_{i}$ is the covariance matrix for state $i$. This joint estimation process ensures that the model parameters are optimally adjusted for the diverse parameter data. The MAP estimation of state $q_{t}$ using the AHSMM is calculated in the following way:
@@
In client quality diagnostics and forecasting, parameters like computation, communication, and data quality influence decision-making differently. The AHSMM effectively integrates these diverse parameters by capturing temporal dependencies and assigning appropriate weights. This enables accurate client quality assessment, improving forecasting and scheduling in dynamic environments.
@@
where $m(i)$ denotes the mean dwell time in state $i$ and $\sigma^{2}(i)$ is the variance of the dwell time in that state, and the term $\rho$ serves as a proportionality constant, which adjusts the influence of the variance on the overall residence time.
This formulation captures the idea that the total time spent in a quality state is influenced not only by the average time spent there but also by the variability of that time. A higher variance indicates greater uncertainty in the duration spent in state $i$, which can lead to longer overall residence times. By incorporating both mean and variance, we obtain a more comprehensive view of the dynamics in quality states. The proportionality constant $\rho$ is defined as:
@@
The reliability function $R(t+e\Delta t)$ represents the probability that the client remains in the current quality state $i$ at time $t+e\Delta t$. Thus, we can get
@@
Based on the above equations, $\bar{E}(i,e)$ can be expressed as:
@@
According to equations (13,14,15), once the client reaches state $i$ and has a dwell time, the trust level (TL) is determined as:
@@
Considering the configuration in SD-FL, the model aggregation within the cluster at server $s$ can be described as follows:
$\boldsymbol{w}_{s,t}=\sum_{i\in U_{s}}\frac{n_{i}}{\sum n_{i}}\boldsymbol{w}_{i,t},$ (17)
where $U_{s}$ represents the set of clients assigned to server $s$, $n_{i}$ is the size of the local dataset for client $i$, and $\boldsymbol{w}_{i,t}$ denotes the local model of client $i$ at training round $t$. This weighted aggregation ensures that clients with larger datasets contribute proportionally more to the cluster model, improving the robustness of the overall learning process.
@@
$\boldsymbol{g}_{s,t}=\sum_{s=1}^{S}\frac{N_{s}}{\sum N_{s}}\boldsymbol{w}_{s,t},$ (18)
where $N_{s}$ corresponds to the total size of the datasets managed by server $s$, and $S$ is the number of servers. This global aggregation step aligns the models from different clusters, ensuring consistency and convergence across the distributed system.
In the subsequent training round, each client utilizes the updated global model $\boldsymbol{g}_{s,t}$ as the initialization for their local model updates. Clients then train their local models using their respective datasets through the gradient descent mechanism, defined as follows:
$\boldsymbol{w}_{i,t+1}=\boldsymbol{w}{i,t}-\eta\nabla F_{i}(\boldsymbol{w}_{i,t}),$ (19)
where $\eta$ is the learning rate, and $\nabla F_{i}(\boldsymbol{w}_{i,t})$ represents the gradient of the local loss function $F_{i}$ with respect to the current model $\boldsymbol{w}_{i,t}$.
This decentralized training mechanism enables clients to collaboratively train a global model while keeping their data local, thereby addressing privacy concerns and minimizing the communication overhead associated with transmitting raw data. The combination of local training, cluster-level aggregation, and global consensus helps achieve a balance between computational efficiency, communication cost, and model accuracy in distributed learning systems.
@@
Client scheduling determines the optimal client-server association matrix $\boldsymbol{d}$ to minimize the global model loss. Each element $d_{ij}$ is binary (1 if client $i$ is assigned to server $j$, 0 otherwise). The configuration of $\boldsymbol{d}$ directly impacts the global loss by influencing data locality, communication overhead, and computational load balancing. Adaptive scheduling dynamically adjusts $\boldsymbol{d}$ to further enhance system performance and ensure efficient training. The global loss function is defined as:
@@
We rely on current trust levels of client $i$ to tackle these challenges, reduce data loss, and guarantee consistent model updates. $\Theta$ denotes the threshold for the trust level of clients involved in training. Reformulating the optimization problem to incorporate parameters that reflect communication link stability will enhance the modeling of the SD-FL system’s conditions.
We also plan to enhance SD-FL system robustness through redundancy strategies, like multiple communication paths or backup servers, to mitigate risks from unreliable links. Dynamically adjusting client-server associations based on real-time assessments will help maintain optimal performance despite trust fluctuations. This approach maximizes resource utilization and minimizes training time, leading to more robust convergence and broader adoption in real-world applications.
@@
By setting the learning rate $\lambda=\frac{1}{L}$, the upper bound of the expected difference $\mathbb{E}\left(F\left(\boldsymbol{g}_{t+1}\right)-F\left(\boldsymbol{g}^{*}\right)\right)$ can be established as follows:
@@
where $D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}$ , $B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi$ , and $\Psi=\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i,s}<\Theta\right)\right)\right)$.